# Lossless data compression at pragmatic rates

### v1

@@ 1.2 Background
The first-order expansion (3) is simply the “rate version” of the well-known error-exponents result that was established in the early works [9, 12] and in the present form by Blahut [4]; see the text [7] for a standard exposition.
@@ 1.3 Compression at pragmatic rates
Strassen: $R_{n}^{*}\approx H(P)+\frac{1}{\sqrt{n}}\sigma(X)Q^{-1}(\epsilon)-\frac{1}{2n}\log n,$ (6)
@@ 2 Preliminaries and auxiliary results
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{Z_{\alpha}}\quad x\in A,$ (9)
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$ All logarithms are taken to base 2.
@@ 3.2 Converse
$N_{0}$ $=\max\Bigg{\{}4.4\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}^{2},\frac{4(1+q+r)^{2}}{p^{2}},\frac{2(1+q+r)}{p(1-\alpha^{*})},N_{1},N_{2}\Bigg{\}},$
$C$ $=\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\hat{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\tilde{\sigma}_{3}^{2}+\hat{\rho_{3}})+1,$
@@
$\sigma_{3}^{2}=\sigma_{3,\alpha^{*}}^{2},\;\;\hat{\sigma_{3}}^{2}=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\tilde{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\rho_{3}}=\sup_{\alpha\in(0,1)}\rho_{3,\alpha},$
$p=(\log e)(1-\alpha^{*})\sigma_{3}^{2},\;\;q=\frac{\log e}{2}(\tilde{\sigma}_{3}^{2}+\hat{\rho}_{3}),\;\;r=19(\log e)(1-\alpha^{*})\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}},$
@@
$P_{\alpha_{n}}^{n}(S_{n}^{c})\leq\frac{1}{2}+\frac{\frac{\rho_{2,\alpha_{n}}}{(1-\alpha_{n})^{3}}}{\frac{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}{(1-\alpha_{n})^{3}}}=\frac{1}{2}+\frac{\rho_{3,\alpha_{n}}}{2\sigma_{3,\alpha_{n}}^{3}\sqrt{n}}\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}\sqrt{n}},$ (23)
where $\sigma_{3,\alpha_{n}}$ and $\rho_{3,\alpha_{n}}$ are as in Lemma 2.6, $\hat{\sigma_{3}}^{2}:=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$ and $\hat{\rho_{3}}:=\sup_{\alpha\in(0,1)}\rho_{3,\alpha}$.
@@
$P^{n}(S_{n})\geq\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}\sqrt{n}}}P^{n}(S).$
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}\sqrt{n}}$, and $\Delta=1$, yields,
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (24)
$\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (25)
as long as $\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}<1$, i.e., for $n>4(\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}+1)^{2}$. Note that $\frac{\hat{\rho_{3}}}{2\hat{\sigma_{3}}^{3}}\geq\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}$ by (23), which justifies (25) and also means that (24) still holds when $n>4(\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1)^{2}$.
@@
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\tilde{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\tilde{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that,
@@
$D(P_{\alpha_{n}}\|P)\leq D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\tilde{\sigma}_{3}^{2}+\hat{\rho}_{3}).$ (27)
@@
$\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}$ $=\Phi^{-1}\Big{(}\frac{1}{2}\Big{)}+\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}\frac{1}{\phi(\Phi^{-1}(x))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(\Phi^{-1}(\frac{1}{2}+\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(2)}\leq\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)},$ (29)
for some $x\in(\frac{1}{2},\frac{1}{2}+\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}})$, where (29) holds as long as:
$n\geq 4.4\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}^{2}>\frac{(\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1)^{2}}{(\Phi(2)-1/2)^{2}}.$
@@
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha^{*}}\|P)-n(\alpha_{n}-\alpha^{*})(\log e)(\alpha^{*}-1)\sigma_{3}^{2}-\frac{n\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\tilde{\sigma}_{3}^{2}+\hat{\rho}_{3})$
$\hskip 43.36243pt-(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}\times(\log e)\sqrt{n}\times\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\hat{\sigma_{3}}^{3}}+1\Big{)}-\frac{1}{2}\log n$
@@
$D(P_{\alpha_{n}}\|P)$ $\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Big{|}\hat{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (34)
$\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2n}\Big{|}\hat{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (35)
$=\delta-\frac{\log n}{2n}-\frac{1+q+r}{n}-\frac{\log e}{2n}\Big{|}\hat{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|},$ (36)
for some $u\in[\alpha^{*},\alpha_{n}]$, where (34) holds because $\sigma_{3,u}^{2}\geq\hat{\sigma_{3}}^{2}$, $1-u\leq 1-\alpha^{*}$ and by the definition of $\hat{\rho_{3}}$; (35) follows from the definition of $\alpha_{n}$; and (36) follows by substituting the expression of $\alpha_{n}$ in (31). Similarly expanding $H(P_{\alpha_{n}})$ around $\alpha^{*}$ and using the expressions for the derivatives of $H(P_{\alpha})$ in Lemma 2.9,
@@
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\tilde{\sigma}_{3}^{2}+1\times\hat{\rho_{3}})$ (38)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2n}(\tilde{\sigma}_{3}^{2}+\hat{\rho_{3}})$ (39)
$=H(P_{\alpha^{*}})-\frac{\alpha^{*}}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{\alpha^{*}(1+q+r)}{1-\alpha^{*}}\frac{1}{n}-\frac{\log e}{2n}(\tilde{\sigma}_{3}^{2}+\hat{\rho_{3}}),$ (40)
where (37) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (38) holds since $\sigma_{3,u}^{2}\leq\tilde{\sigma_{3}}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho_{3}}$; (39) follows from the definition of $\alpha_{n}$; and (40) follows by substituting the expression of $\alpha_{n}$ in (31).
@@
$-\Bigg{\{}\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\hat{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\tilde{\sigma}_{3}^{2}+\hat{\rho_{3}})\Bigg{\}}\times\frac{1}{n},$

### v2

@@ Abstract
00footnotetext: This work was supported in part by the EPSRC-funded INFORMED-AI project EP/Y028732/1.
@@ 1.2 Background
The first-order asymptotic relation (3) is simply the “rate version” of the well-known error-exponents result that was established in the early works [9, 12] and in the present form by Blahut [4]. Indeed, $H(P_{\alpha^{*}})$ can be expressed as the inverse of the error-exponent function,
$\Delta(R):=\inf_{P^{\prime}:H(P^{\prime})\geq R}D(P^{\prime}\|P),$
via,
$H(P_{\alpha^{*}})=\Delta^{-1}(\delta)=\sup_{P^{\prime}:D(P^{\prime}\|P)\leq\delta}H(P^{\prime}),$
where $D(P^{\prime}\|P)$ is the relative entropy between $P^{\prime}$ and $P$.
@@ 1.3 Compression at pragmatic rates
Strassen: $R_{n}^{*}\approx H(P)+\frac{1}{\sqrt{n}}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2n}\log n,$ (6)
@@ 2 Preliminaries and auxiliary results
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{Z_{\alpha}},\quad x\in A,$ (9)
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, ‘$\log$’ denotes the logarithm taken to base 2, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$
@@ 3.2 Converse
$N_{0}$ $=\max\Bigg{\{}4.4\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}^{2},\frac{4(1+q+r)^{2}}{p^{2}},\frac{2(1+q+r)}{p(1-\alpha^{*})},N_{1},N_{2}\Bigg{\}},$
$C$ $=\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})+1,$
@@
$\sigma_{3}^{2}=\sigma_{3,\alpha^{*}}^{2},\;\;\tilde{\sigma_{3}}^{2}=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\rho_{3}}=\sup_{\alpha\in(0,1)}\rho_{3,\alpha},$
$p=(\log e)(1-\alpha^{*})\sigma_{3}^{2},\;\;q=\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),\;\;r=19(\log e)(1-\alpha^{*})\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}},$
@@
$P_{\alpha_{n}}^{n}(S_{n}^{c})\leq\frac{1}{2}+\frac{\frac{\rho_{2,\alpha_{n}}}{(1-\alpha_{n})^{3}}}{\frac{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}{(1-\alpha_{n})^{3}}}=\frac{1}{2}+\frac{\rho_{3,\alpha_{n}}}{2\sigma_{3,\alpha_{n}}^{3}\sqrt{n}}\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}},$ (23)
where $\sigma_{3,\alpha_{n}}$ and $\rho_{3,\alpha_{n}}$ are as in Lemma 2.6, $\tilde{\sigma_{3}}^{2}:=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$ and $\hat{\rho_{3}}:=\sup_{\alpha\in(0,1)}\rho_{3,\alpha}$.
@@
$P^{n}(S_{n})\geq\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}}P^{n}(S).$
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}$, and $\Delta=1$, yields,
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (24)
$\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (25)
as long as $\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}<1$, i.e., for $n>4(\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}+1)^{2}$. Note that $\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}}\geq\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}$ by (23), which justifies (25) and also means that (24) still holds when $n>4(\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1)^{2}$.
@@
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\hat{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that,
@@
$D(P_{\alpha_{n}}\|P)\leq D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}).$ (27)
@@
$\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}$ $=\Phi^{-1}\Big{(}\frac{1}{2}\Big{)}+\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}\frac{1}{\phi(\Phi^{-1}(x))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(\Phi^{-1}(\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(2)}\leq\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)},$ (29)
for some $x\in(\frac{1}{2},\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}})$, where (29) holds as long as:
$n\geq 4.4\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}^{2}>\frac{(\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1)^{2}}{(\Phi(2)-1/2)^{2}}.$
@@
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha^{*}}\|P)-n(\alpha_{n}-\alpha^{*})(\log e)(\alpha^{*}-1)\sigma_{3}^{2}-\frac{n\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})$
$\hskip 43.36243pt-(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}\times(\log e)\sqrt{n}\times\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}-\frac{1}{2}\log n$
@@
$D(P_{\alpha_{n}}\|P)$ $\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (34)
$\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2n}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (35)
$=\delta-\frac{\log n}{2n}-\frac{1+q+r}{n}-\frac{\log e}{2n}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|},$ (36)
where (34) holds because the right-hand side is a lower bound to (26) since $\sigma_{3,u}^{2}\geq\tilde{\sigma_{3}}^{2}$, $1-u\leq 1-\alpha^{*}$ and by the definition of $\hat{\rho_{3}}$; (35) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (36) follows by substituting the expression of $\alpha_{n}$ in (31). Similarly expanding $H(P_{\alpha_{n}})$ around $\alpha^{*}$ and using the expressions for the derivatives of $H(P_{\alpha})$ in Lemma 2.9,
@@
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+1\times\hat{\rho_{3}})$ (38)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})$ (39)
$=H(P_{\alpha^{*}})-\frac{\alpha^{*}}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{\alpha^{*}(1+q+r)}{1-\alpha^{*}}\frac{1}{n}-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}}),$ (40)
where (37) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (38) holds since $\sigma_{3,u}^{2}\leq\hat{\sigma_{3}}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho_{3}}$; (39) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (40) follows by substituting the expression of $\alpha_{n}$ in (31).
@@
$-\Bigg{\{}\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})\Bigg{\}}\times\frac{1}{n},$